OpenStudy (blackops2_143):

What is the solution to the equation? 8x + 4 = –20 A. x = –3 B. x = –2 C. x = 2 D. x = 3

alones (alones):

Let's solve your equation step-by-step. 8x+4=−20 Step 1: Subtract 4 from both sides. 8x+4−4=−20−4 8x=−24 Step 2: Divide both sides by 8. 8x8=−248 x=−3

OpenStudy (blackops2_143):

ok